Use the following go show representation of square numbers
Square grid
Isometric dotty paper
Square dotty paper
1
Expert's answer
2021-07-13T07:04:05-0400

1) Square grid


3 * 3 = 9

2) Isometric dotty paper

Using 3d we can draw a cube on the isometric dotty paper. A cube contains square sides as shown. Hence 3×3=9


3)

Square dotty paper



10×10=100